# Closing the Confidence-Faithfulness Gap in Large Language Models

### 26 Mar 2026 v1

@@ 1 Introduction
A growing body of mechanistic-interpretability research has shown that high-level semantic and behavioral properties are encoded as linear directions in the residual stream. Linear probes recover truth and falsehood from internal activations (Burns et al., 2024; Marks and Tegmark, 2024; Azaria and Mitchell, 2023), and steering vectors along these directions causally shift model behavior at inference time for truthfulness (Li et al., 2023), broad behavioral traits (Zou et al., 2025; Turner et al., 2024; Rimsky et al., 2024), and refusal (Arditi et al., 2024). Recent work has begun extending this lens to verbalized confidence. Kumaran et al. (2026) show that verbal confidence is cached at answer-adjacent positions and retrieved later. Seo et al. (2026) identify “answer-independence” as a driver of overconfidence and propose a fine-tuning fix. These studies establish that verbalized confidence has a nontrivial internal presence, yet a core question remains unanswered: what is the geometric relationship between the model’s internal accuracy signal and its verbalized confidence, and can that relationship be leveraged to improve calibration?
@@ 2.2 Contrastive Activation Steering
We conduct experiments across three model families: Llama-3.1-8B (Grattafiori and et al, 2024), Qwen2.5-7B (Qwen et al., 2025), and Mistral-7B-v0.3 (Jiang et al., 2023). For each family, we analyze both the base (pretrained) model and its corresponding instruction-tuned (instruct) variant.
@@ 3.2 Verbalized Confidence Is Linearly Separable
Using activations from the pure confidence prompt, we project onto the first principal component and color each point by whether the model expressed high or low confidence. Figure 2 reveals clear linear separability between high- and low-confidence activations in later layers, suggesting that the model progressively constructs a linearly separable representation of its own confidence. To quantify this effect, we train linear ridge regression and report train and test time $E^{2}$ in Figure 3 a and b. The natural next question is whether they share the same direction or dissociated, which would explain verbalized miscalibration.
@@ 3.6 Cross-Dataset Generalization
Table 2: MATH-derived steering vectors transfer across datasets. Mean verbalized confidence (%) under varying steering magnitudes, using a CAA vector trained only on MATH at layer 21 of Qwen2.5-7B, layer 24 of Llama-3.1-8B, and layer 27 of Mistral-7B-v0.3. Takeaway: The confidence direction is domain-general, not an artifact of mathematical notation or problem format. Results increase (→) across columns.
Negative Steering Positive Steering Target Dataset $\alpha{=}{-}0.75$ 0.50{-}0.50 0.25{-}0.25 0 $0.25$ $0.50$ $0.75$ Llama-3.1-8B-Instruct    MMLU 3.6 6.4 13.8 18.7 34.3 56.6 76.2    TriviaQA 6.0 7.2 15.6 24.2 30.4 57.5 76.5    TruthfulQA 5.6 9.7 19.5 21.9 35.7 57.7 75.7 Mistral-7B-v0.3-Instruct    MMLU 8.1 18.5 22.6 33.8 50.5 74.7 86.4    TriviaQA 10.8 18.5 26.1 41.0 54.6 75.3 85.0    TruthfulQA 12.9 20.9 27.5 43.2 58.6 73.5 85.4 Qwen2.5-7B-Instruct    MMLU 58.3 59.1 64.2 64.0 67.1 66.3 68.8    TriviaQA 52.7 54.9 57.9 59.2 60.6 61.8 64.3    TruthfulQA 52.2 54.3 59.1 59.6 59.9 61.6 62.9
@@ 3.7 Base-to-Instruct Transfer
Neg. Steering Pos. Steering Model $\alpha{=}{-}0.75$ 0.50{-}0.50 0.25{-}0.25 0 $0.25$ $0.50$ $0.75$ Qwen2.5-7B-Inst. 38.6 75.5 86.5 88.4 90.3 94.1 98.6 Llama-3.1-8B-Inst. 11.4 39.2 85.8 87.8 92.3 94.4 95.8 Mistral-7B-Inst.-v0.3 83.3 88.1 90.4 93.6 92.6 92.6 94.6
Table 3 shows that base-model-derived steering vectors successfully modulate instruct model confidence across all three model families, albeit the weaker signal on Mistral. This result has two implications. First, the linear confidence direction identified in base models is not completely eliminated by post-training; it persists in the instruct model’s residual stream in a geometrically compatible form and remains stronger in some models than others. Second, it is possible that the overconfidence exhibited by instruct models is not a consequence of losing the confidence signal entirely, but rather of the generation process failing to read it out faithfully. Thus, steering could provide a direct mechanism to restore confidence control in instruct-tuned models.
@@ 3.8 Adaptive Two-Stage Steering for Verbalized Calibration Improvement
Model Confidence Source ECE $\downarrow$ Brier $\downarrow$ MAE $\downarrow$ Llama-3.1-8B-Inst. Logit baseline 68.4 50.5 68.5 Verbalized 14.9 8.4 20.2 SteerConf 46.3 40.6 53.6 Verbalized (steered) 3.7 2.4 10.5 Mistral-7B-Inst.-v0.3 Logit baseline 73.5 57.5 73.5 Verbalized 35.1 15.9 40.0 SteerConf 24.3 20.1 29.1 Verbalized (steered) 3.3 2.1 9.3 Qwen2.5-7B-Inst. Logit baseline 75.0 58.3 74.7 Verbalized 36.0 16.7 36.9 SteerConf 19.5 25.9 41.4 Verbalized (steered) 10.7 3.5 11.8

### 01 Apr 2026 v2

@@ 1 Introduction
A growing body of mechanistic-interpretability research has shown that high-level semantic and behavioral properties are encoded as linear directions in the residual stream. Linear probes recover truth and falsehood from internal activations (Burns et al., 2024; Marks and Tegmark, 2024; Azaria and Mitchell, 2023), and steering vectors along these directions causally shift model behavior at inference time for truthfulness (Li et al., 2023), broad behavioral traits (Zou et al., 2025; Turner et al., 2024; Rimsky et al., 2024), and refusal (Arditi et al., 2024). Simultaneous work has begun extending this lens to verbalized confidence. Kumaran et al. (2026) show that verbal confidence is cached at answer-adjacent positions and retrieved later. Seo et al. (2026) identify “answer-independence” as a driver of overconfidence and propose a fine-tuning fix. These studies establish that verbalized confidence has a nontrivial internal presence, yet a core question remains unanswered: what is the geometric relationship between the model’s internal accuracy signal and its verbalized confidence, and can that relationship be leveraged to improve calibration?
@@ 2.2 Contrastive Activation Steering
We conduct experiments across three model families: Llama-3.1-8B (Grattafiori and et al, 2024), Qwen2.5-7B (Qwen et al., 2025), and Mistral-7B-v0.1 (Jiang et al., 2023). For each family, we analyze both the base (pretrained) model and its corresponding instruction-tuned (instruct) variant.
@@ 3.2 Verbalized Confidence Is Linearly Separable
Using activations from the pure confidence prompt, we project onto the first principal component and color each point by whether the model expressed high or low confidence. Figure 2 reveals clear linear separability between high- and low-confidence activations in later layers, suggesting that the model progressively constructs a linearly separable representation of its own confidence. To quantify this effect, we train linear ridge regression and report train and test time $R^{2}$ in Figure 3 a and b. The natural next question is whether they share the same direction or dissociated, which would explain verbalized miscalibration.
@@ 3.6 Cross-Dataset Generalization
Table 2: MATH-derived steering vectors transfer across datasets. Mean verbalized confidence (%) under varying steering magnitudes, using a CAA vector trained only on MATH at layer 21 of Qwen2.5-7B, layer 24 of Llama-3.1-8B, and layer 27 of Mistral-7B-v0.1. Takeaway: The confidence direction is domain-general, not an artifact of mathematical notation or problem format. Results increase (→) across columns.
Negative Steering Positive Steering Target Dataset $\alpha{=}{-}0.75$ 0.50{-}0.50 0.25{-}0.25 0 $0.25$ $0.50$ $0.75$ Llama-3.1-8B    MMLU 3.6 6.4 13.8 18.7 34.3 56.6 76.2    TriviaQA 6.0 7.2 15.6 24.2 30.4 57.5 76.5    TruthfulQA 5.6 9.7 19.5 21.9 35.7 57.7 75.7 Mistral-7B-v0.1    MMLU 8.1 18.5 22.6 33.8 50.5 74.7 86.4    TriviaQA 10.8 18.5 26.1 41.0 54.6 75.3 85.0    TruthfulQA 12.9 20.9 27.5 43.2 58.6 73.5 85.4 Qwen2.5-7B    MMLU 58.3 59.1 64.2 64.0 67.1 66.3 68.8    TriviaQA 52.7 54.9 57.9 59.2 60.6 61.8 64.3    TruthfulQA 52.2 54.3 59.1 59.6 59.9 61.6 62.9
@@ 3.7 Base-to-Instruct Transfer
Neg. Steering Pos. Steering Model $\alpha{=}{-}0.75$ 0.50{-}0.50 0.25{-}0.25 0 $0.25$ $0.50$ $0.75$ Qwen2.5-7B-Inst. 38.6 75.5 86.5 88.4 90.3 94.1 98.6 Llama-3.1-8B-Inst. 11.4 39.2 85.8 87.8 92.3 94.4 95.8 Mistral-7B-Inst.-v0.1 37.6 46.7 50.0 63.2 80.5 86.6 90.2
Table 3 shows that base-model-derived steering vectors successfully modulate instruct model confidence across all three model families. This result has two implications. First, the linear confidence direction identified in base models is not completely eliminated by post-training; it persists in the instruct model’s residual stream in a geometrically compatible form and remains stronger in some models than others. Second, it is possible that the overconfidence exhibited by instruct models is not a consequence of losing the confidence signal entirely, but rather of the generation process failing to read it out faithfully. Thus, steering could provide a direct mechanism to restore confidence control in instruct-tuned models.
@@ 3.8 Adaptive Two-Stage Steering for Verbalized Calibration Improvement
Model Confidence Source ECE $\downarrow$ Brier $\downarrow$ MAE $\downarrow$ Llama-3.1-8B Logit baseline 68.4 50.5 68.5 Verbalized 14.9 8.4 20.2 SteerConf 46.3 40.6 53.6 Verbalized (steered) 3.7 2.4 10.5 Mistral-7B–v0.1 Logit baseline 73.5 57.5 73.5 Verbalized 35.1 15.9 40.0 SteerConf 24.3 20.1 29.1 Verbalized (steered) 3.3 2.1 9.3 Qwen2.5-7B Logit baseline 75.0 58.3 74.7 Verbalized 36.0 16.7 36.9 SteerConf 19.5 25.9 41.4 Verbalized (steered) 10.7 3.5 11.8